OpenStudy (anonymous):

sqrt x+20-x=0 being x+20 is all that is under the root.

OpenStudy (anonymous):

\[\sqrt{x+20} - x =0\]

OpenStudy (anonymous):

\[\sqrt{x+20} = x\]

OpenStudy (anonymous):

sqrtx+20=X square both sides x+20=x^2 (x-5)(x+4)

OpenStudy (anonymous):

harkirat has the equation correct

OpenStudy (anonymous):

square both sides x + 20 = x^2 so u get quadratic equation x^2 - x - 20 = 0 now factorise it and get the two roots/values of x

OpenStudy (anonymous):

accordingly x=5,x=-4

OpenStudy (anonymous):

The first equation harkitar has is right

OpenStudy (anonymous):

x^2 - 5x + 4x - 20 = 0 x(x -5) +4(x -5) = 0 (x+4)(x-5) = 0 so x = -4 , 5

OpenStudy (anonymous):

pls click GOOD ANSWER button if u r satisfied with the answer (: }))